OpenStudy (anonymous):

ABCD is a rectangle. find the length of each diagonal AC=c/3 BD=4-c

OpenStudy (freckles):

I like drawing I would start with a drawing but you will definitely just really need Pythagorean theorem

OpenStudy (freckles):

oops actually those aren't the side measurement given

OpenStudy (freckles):

those are the diagonals

OpenStudy (freckles):

this is easier than I thought

OpenStudy (freckles):

what do you know about the 2 diagonals of a rectangle? like about their relationship?

OpenStudy (anonymous):

im taking a guess but do they have the same measurements

OpenStudy (freckles):

yep yep yep

OpenStudy (freckles):

c/3=4-c solve this for c

OpenStudy (anonymous):

it would be 1 right?

OpenStudy (freckles):

well if we replace c with 1 we get 1/3=4-1 and 1/3 isn't 3

OpenStudy (freckles):

try multiplying both sides by 3 so you can get rid of that annoying fraction c=12-3c see if you can solve from here

OpenStudy (triciaal):